OpenStudy (anonymous):

Calculate the molality of each of the following solutions. 1. 93.2 g of NaCl in 583 g of H2O

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

but first you need to convert the mass of solute to moles.

OpenStudy (anonymous):

so do i need to get 93.2 into liters and 583 g of h2o into liters?

OpenStudy (aaronq):

ahh no, i messed up. i wrote the molarity formula. it should be: \(molality=\dfrac{moles~of~solute}{kilograms~of~solvent}\)

OpenStudy (anonymous):

I'm confused?

OpenStudy (aaronq):

i wrote the wrong formula before. This one ^ is right. Do you remember how to find the moles of something? \(moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (aaronq):

find the moles of the solute first.

OpenStudy (anonymous):

wat?

OpenStudy (aaronq):

to find the molality of the solution, you need to know the moles of the solute. find the moles of the solute

OpenStudy (anonymous):

so i divide 93.2 over moles and 583 over moles

OpenStudy (aaronq):

no, not yet. you need to know the moles of solute, you only know the mass. Divide the mass of the solute (the salt) by the molar mass to get the moles.

OpenStudy (anonymous):

so divide 58.4425 g/mol

OpenStudy (anonymous):

by 22.98

OpenStudy (aaronq):

93.2 g/58.4425 g/mol = moles of NaCl then \(molality=\dfrac{moles~of ~NaCl}{kilograms~of~water}\) notice that you don't need to find the moles of water, you use the mass

OpenStudy (anonymous):

so does 1.5947 go on top or bottom?

OpenStudy (aaronq):

look at the formula, it goes on the top.

OpenStudy (anonymous):

ok and then 93.2 converts to kilo?

OpenStudy (aaronq):

take 583 g and convert it to kg.

OpenStudy (anonymous):

it takes 583 equals=0.0583?

OpenStudy (aaronq):

583/1000=0.583

OpenStudy (anonymous):

then?

OpenStudy (aaronq):

so now plug these 2 values into the molality formula.

OpenStudy (anonymous):

so 1.5947/0.583?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

so its 2.735?

OpenStudy (aaronq):

yep! thats it

OpenStudy (anonymous):

but wats the unit

OpenStudy (aaronq):

"m" (for molality) or "mol/kg of solvent"

OpenStudy (anonymous):

is that the final answer?

OpenStudy (aaronq):

yes, it is.

OpenStudy (anonymous):

ok i have a final question and thank you so much for the help message me ur address i wanna mail u something to make it up to u

OpenStudy (aaronq):

no it's all good! no problem post it!

OpenStudy (anonymous):

ok thanks ur really nice